OpenStudy (anonymous):

The perimeter of a rectangle is 22ft , and the area of the rectangle is 24ft^2 . Find the dimensions of the rectangle.

OpenStudy (anonymous):

perimeter is the sum of all the sides area of a rectangle is A= L * H where L - length and H - height

OpenStudy (anonymous):

perimeter of a rectangle can be written as P= 2L +2 H

OpenStudy (anonymous):

2l + 2w = 22 ft and w*l = 24 ft^2 from first equation: w=(22-2l)/2 = 11 - l substitute in eq, 2 (11-l)*l=24 ft^2 11*l - l^2 = 24 solve for l, l = 3 or 8 if l=3, w= 8 if l=8, w=3

OpenStudy (anonymous):

thank you